### v1

@@ 1 Introduction
We then propose alternative sets of assumptions under which the DiD estimand becomes informative about specific causal effects. These assumptions leverage information on the trend of the potential outcome in the absence of treatment and on the nature of the interference to partially identify and disentangle the total effect on the treated and the indirect effect on the control separately. We illustrate our results by revisiting the seminal paper on minimum wages and employment by Card and Krueger, (1994). We discuss why the no-interference assumption embedded in SUTVA may be violated in their application, how such violations impact the interpretation of their DiD estimates, and which additional assumptions are needed for their original conclusions to remain valid.
@@ 2 Difference-in-Differences under interference
This section introduces the canonical DiD research design and examines what the DiD estimand identifies in the presence of interference. For simplicity, we consider the two-groups, two-periods case. Similar arguments can be employed for the multiple-groups, multiple-periods case.
@@ 2.1 Notation and set-up
Let $Y$ denote the outcome of interest. We postulate the existence of $2^{N\times 2}$ potential outcomes for unit $i$ at time $t$, $Y_{it}(\boldsymbol{W})$, where $\boldsymbol{W}$ is the $N\times 2$ treatment assignment matrix containing the potential treatment assignment for all units in all periods. We can decompose this matrix into $N\times 1$ vectors of treatment assignment, one in each period, $Y_{it}(\boldsymbol{W_{0}},\boldsymbol{W_{1}})$, where $\boldsymbol{W_{t}}$ denotes the treatment assignment vector at time $t$. By definition, no unit is treated in period $t=0$, so we can write $Y_{it}(\boldsymbol{0},\boldsymbol{W_{1}})$.
@@
Assumption 1 states that outcomes depend only on past and current treatments, but not on future ones. This assumption implies that the treatment cannot influence outcomes measured before its implementation. Therefore, we can express the potential outcomes in the pre-treatment period as a function of the contemporaneous treatment vector only, $Y_{i0}(\boldsymbol{W_{0}})=Y_{i0}(\boldsymbol{0})$. Since no unit is treated at $t=0$ by design, the potential outcomes at $t=1$ can be indexed exclusively by the treatment assignments at that time, $\boldsymbol{W_{1}}$. Under interference and the no-anticipation assumption, each unit is characterized by $2^{N}$ potential outcomes in each period.
@@
Assumption 2, also known as consistency (Rubin,, 1980), links the observed outcome for each unit to a single well-defined potential outcome.
@@ 2.2 Canonical DiD under the no-interference assumption
Assumption 3 states that potential outcomes do not depend on other units’ treatment. It rules out both interference between and within groups. This assumption is strong and is likely to be violated in many scenarios. Under Assumption 3, it suffices to index the potential outcomes only with the unit’s own treatment, which is why it is embedded in the notation used by researchers, even if it is not explicitly stated (Arkhangelsky and Imbens,, 2024).
@@
It is clear that, without Assumption 3, this estimand is ill-defined. Estimand 1 captures the expected effect of treating a single unit from the treatment group. The DiD research design is typically applied to observational settings where the treatment is confounded with unobservable characteristics. Therefore, the identification of the ATT estimand relies on assumptions on the evolution of these unobservable confounders, given Assumptions 1 and 3. The main identification assumption in the canonical DiD framework is the Parallel Trends assumption.
@@
where Assumptions 1, 2, and 3 are used in the second equality to postulate the potential outcomes and link them to the observed outcomes, and the third equality follows from Assumption 4.
@@ 2.3 DiD under partial interference
Assumption 5 rules out interference between treated and control units, similar to the SUTNVA assumption in Forastiere et al., (2021). This assumption is a weaker version of the no-interference assumption embedded in Assumption 3. When Assumption 3 holds, Assumption 5 also holds. However, the opposite is not true, as even when Assumption 5 holds, there might be spillover among the treated units. The immediate consequence of this relaxation is that the main target estimand in the DiD setting, the ATT, is no longer well-defined, since potential outcomes may now also depend on the treatment assignment of other units in the same group. Under partial interference, the following causal estimand is well-defined:
@@
where assumptions 1, 2, and 5 are used in the second equality to postulate the potential outcomes and link them to the observed outcomes, and the third equality follows from assumption 6.
Proposition 2 states that the DiD estimand has a causal interpretation when units interfere within groups, provided that there is no interference between groups. However, it no longer identifies the ATT, which is not well-defined without Assumption 3, but the TATT-pi. This distinction changes the interpretation of the DiD estimand. Now, it incorporates both the average direct treatment effect and the average spillover effect on the treated. Thus, it should be interpreted as the expected effect when treating the whole group, rather than a single unit. The gap between these two interpretations becomes especially important when scaling up a policy, as changing the composition and size of the treated group may lead to different results. Note that this result is derived under Assumption 5, which does not impose any constraint on the interference structure in the treatment group.
@@ 2.4 DiD under unknown interference
Assumption 7 states that, in the absence of the intervention, the difference in expected outcomes between the two groups would have remained constant over time. Therefore, the interpretation of this parallel trends assumption is equivalent to the parallel trends assumption under no-interference (Assumption 4) and the parallel trends assumption under partial interference (Assumption 6).
Observed Outcome Potential Outcome
$\mathbb{E}[Y_{i1}\mid G_{i}=1]$ $\mathbb{E}[Y_{i1}(1,\boldsymbol{1},\boldsymbol{0})\mid G_{i}=1]$
$\mathbb{E}[Y_{i0}\mid G_{i}=1]$ $\mathbb{E}[Y_{i0}(0,\boldsymbol{0},\boldsymbol{0})\mid G_{i}=1]$
$\mathbb{E}[Y_{i1}\mid G_{i}=0]$ $\mathbb{E}[Y_{i1}(0,\boldsymbol{0},\boldsymbol{1})\mid G_{i}=0]$
$\mathbb{E}[Y_{i0}\mid G_{i}=0]$ $\mathbb{E}[Y_{i0}(0,\boldsymbol{0},\boldsymbol{0})\mid G_{i}=0]$
Table 1: Correspondence between the observed outcomes and the potential outcomes.
@@ 3 Alternative identification strategies under unknown interference
Constant average potential outcomes in the absence of treatment:
@@
where Assumption 1 and 2 are used in the first equality to postulate the potential outcomes and link them to the observed outcomes, and the second equality follows from Assumption 8. Similar arguments can be used to identify $\tau^{0}$.
Assumption 8 is strong, yet it illustrates an important insight: in the presence of interference, even if the potential outcomes $Y_{1}(0,0,0)$ were observed, we cannot compare treatment and control groups to identify causal effects, as both groups are affected (potentially differently) by the treatment. A more general version of this assumption is given by:
@@
$\mathbb{E}[Y_{it+1}(0,\boldsymbol{0},\boldsymbol{0})\mid G_{i}=g]=\mathbb{E}[Y_{it}(0,\boldsymbol{0},\boldsymbol{0})\mid G_{i}=g]\pm k,$
@@ 4 Application: revisiting Card and Krueger, (1994)
However, a substantial body of literature in economics demonstrates that increases in minimum wages generate spillover effects (Grossman,, 1983; Cengiz et al.,, 2019; Caires,, 2024). Part of this literature focuses on spatial spillovers and demonstrates that the increase of minimum wages may also affect bordering regions in various ways (Kuehn,, 2016; Shirley,, 2018; McKinnish,, 2017; Jardim et al.,, 2022; Jha et al.,, 2024), with commuting being one of the main drivers of spillovers. In the presence of general equilibrium effects, it is reasonable to think that the increase in minimum wage in New Jersey also affected fast food restaurants in eastern Pennsylvania, thereby violating the no-interference assumption. The spillovers could arise from multiple channels and operate in different directions. For instance, in the context of cross-border commuting, some Pennsylvania workers may have sought employment in New Jersey to benefit from higher wages, reducing labor supply in Pennsylvania and exerting downward pressure on restaurant employment there. This mechanism could explain the observed decline in Pennsylvania employment, which drives the positive treatment effect for New Jersey.
If no-interference does not hold, it is no longer possible to claim that the increase in the minimum wage led to an increase in employment. Instead, conditional on the parallel trends assumption (Assumption 7) holding, we can only determine the differential effect. That is, we could conclude that the total average effect of the raise in minimum wage on employment in New Jersey was 2.75 FTE employees larger than the average spillover effect in Pennsylvania. However, we cannot determine the effect on employment in any of the states. In fact, it is possible that the rise in minimum wage decreased employment in both states, with a larger decrease in Pennsylvania, resulting in a positive difference but contradicting the original claim in Card and Krueger, (1994).

### v2

@@ 1 Introduction
We then propose alternative sets of assumptions under which the DiD estimand becomes informative about specific causal effects. These assumptions leverage information on the trend of the potential outcome in the absence of treatment and on the nature of the interference to partially identify and disentangle the total effect on the treated and the indirect effect on the control separately. We illustrate our results by revisiting the seminal paper on minimum wages and employment by Card and Krueger, (1994). We discuss why the no-interference assumption embedded in SUTVA may be violated in their application, how such violations impact the interpretation of their DiD estimates, and which additional assumptions are needed to maintain the validity of their original conclusions. We also relate our framework to some recent empirical studies (Braghieri et al.,, 2022; Fenizia and Saggio,, 2024; Chen et al.,, 2025; Truffa and Wong,, 2025), showing how our formalization clarifies the interpretation of their DiD estimands and provides a structured foundation for arguments that are often made informally.
@@ 2 Difference-in-Differences under interference
This section introduces the canonical DiD research design and examines what the DiD estimand identifies in the presence of interference. For simplicity, we consider the two-group, two-period case. Similar arguments can be employed for the multiple-group, multiple-period case.
@@ 2.1 Notation and set-up
Let $Y$ denote the outcome of interest. We postulate the existence of $2^{N\times 2}$ potential outcomes for unit $i$ at time $t$, $Y_{it}(\boldsymbol{W})$, where $\boldsymbol{W}$ is the $N\times 2$ treatment assignment matrix containing the potential treatment assignment for all units in all periods. We can decompose this matrix into $N\times 1$ vectors of treatment assignment, one in each period, $Y_{it}(\boldsymbol{W_{0}},\boldsymbol{W_{1}})$, where $\boldsymbol{W_{t}}$ denotes the treatment assignment vector at time $t$. In the specific setting we consider, no unit is treated in period $t=0$, so we can write $Y_{it}(\boldsymbol{0},\boldsymbol{W_{1}})$.
@@
Assumption 1 states that outcomes do not depend on future treatments. This assumption implies that treatments cannot affect outcomes measured before their implementation.
Therefore, we can express the potential outcomes in the pre-treatment period as a function of the contemporaneous treatment vector only, $Y_{i0}(\boldsymbol{W_{0}})=Y_{i0}(\boldsymbol{0})$. Since no unit is treated at $t=0$ by design, the potential outcomes at $t=1$ can be indexed exclusively by the treatment assignments at that time, $\boldsymbol{W_{1}}$. In the two-group two-period case, under interference and the no-anticipation assumption, each unit is characterized by $2^{N}$ potential outcomes in each period.
@@
Assumption 2, also known as consistency (VanderWeele and Hernán,, 2013), links the observed outcome for each unit to a single well-defined potential outcome.
@@ 2.2 Canonical DiD under the no-interference assumption
Assumption 3 states that potential outcomes of unit $i$ do not depend on other units’ treatment. It rules out both interference between and within groups. This assumption is strong and is likely to be violated in many scenarios. Under Assumption 3, it suffices to index the potential outcomes only with the unit’s own treatment, which is why it is embedded in the notation used by researchers, even if it is not explicitly stated (Arkhangelsky and Imbens,, 2024).
@@
It is clear that, without Assumption 3, this estimand is ill-defined. Estimand 1 captures the expected effect of treating a single unit drawn from the treatment group. The DiD research design is typically applied to observational settings where the treatment is confounded with unobservable characteristics. Therefore, the identification of the ATT estimand relies on assumptions on the evolution of these unobservable confounders, given Assumptions 1 and 3. The main identification assumption in the canonical DiD framework is the Parallel Trends assumption.
@@
where Assumptions 1, 2, and 3 are used in the second equality to link observed outcomes to potential outcomes, and the third equality follows from Assumption 4.
@@ 2.3 DiD under partial interference
Assumption 5 rules out interference between treated and control units, similar to the SUTNVA assumption in Forastiere et al., (2021). This assumption is a weaker version of the no-interference assumption embedded in Assumption 3. When Assumption 3 holds, Assumption 5 also holds. However, the opposite is not true, because even when Assumption 5 holds, there might be spillovers among the treated units. The immediate consequence of this relaxation is that the main target estimand in the DiD setting, the ATT, is no longer well-defined, since potential outcomes may now also depend on the treatment assignment of other units in the same group. Under partial interference, the following causal estimand is well-defined:
@@
where assumptions 1, 2, and 5 are used in the second equality to link observed outcomes to potential outcomes, and the third equality follows from assumption 6.
Proposition 2 states that the DiD estimand has a causal interpretation when units interfere within groups, provided that there is no interference between groups. However, it no longer identifies the ATT, which is not well-defined without Assumption 3, but the TATT-pi. This distinction changes the interpretation of the DiD estimand. Now, it incorporates both the average direct treatment effect and the average spillover effect on the treated. Thus, it should be interpreted as the expected effect when treating the whole group, rather than a single unit. The gap between these two interpretations becomes especially important when scaling up a policy, as changing the composition and size of the treated group may lead to different results. Note that this result is derived under Assumption 5, which does not impose any constraint on the interference structure within the treatment group.
@@ 2.4 DiD under unknown interference
Assumption 7 states that, in the absence of the intervention, the difference in expected outcomes between the two groups would have remained constant over time. Therefore, the interpretation of this parallel trends assumption is equivalent to the parallel trends assumption under no-interference (Assumption 4) and the parallel trends assumption under partial interference (Assumption 6) [^2] In this note, we consider estimands and assumptions whose analogs under SUTVA are well-defined. It is still possible to explore similar parallel trends assumption on other observed potential outcomes under unknown interference, like $Y_{it}(1,\boldsymbol{1},\boldsymbol{0})$ or $Y_{it}(0,\boldsymbol{0},\boldsymbol{1})$, under which identifications of estimands that are well defined only under interference, like the average spillover effects for the treated, could follow..
@@ 3 Alternative identification strategies under unknown interference
Constant average potential outcomes in the absence of treatment, $g\in\{0,1\}$:
@@
where Assumption 1 and 2 are used in the first equality to link the observed outcomes to potential outcomes, and the second equality follows from Assumption 8. Similar arguments can be used to identify $\tau^{0}$.
Assumption 8 is strong, yet it illustrates an important insight: in the presence of interference, even if the potential outcomes $Y_{1}(0,0,0)$ were observed, we could not compare treatment and control groups to identify causal effects, as both groups are affected (potentially differently) by the treatment. A more general version of this assumption is given by:
@@
$\mathbb{E}[Y_{it+1}(0,\boldsymbol{0},\boldsymbol{0})\mid G_{i}=g]-\mathbb{E}[Y_{it}(0,\boldsymbol{0},\boldsymbol{0})\mid G_{i}=g]\in[-k,k],$
@@ 4 Application: revisiting Card and Krueger, (1994)
However, a substantial body of literature in economics demonstrates that increases in minimum wages generate spillover effects (e.g., Grossman, (1983); Cengiz et al., (2019); Caires, (2024)). Part of this literature focuses on spatial spillovers and demonstrates that the increase of minimum wages may also affect bordering regions in various ways (e.g., Kuehn, (2016); Shirley, (2018); McKinnish, (2017); Jardim et al., (2022); Jha et al., (2024)), with commuting being one of the main drivers of spillovers. In the presence of general equilibrium effects, it is reasonable to think that the increase in minimum wage in New Jersey also affected fast food restaurants in eastern Pennsylvania, thereby violating the no-interference assumption. The spillovers could arise from multiple channels and operate in different directions. For instance, in the context of cross-border commuting, some Pennsylvania workers may have sought employment in New Jersey to benefit from higher wages, reducing labor supply in Pennsylvania and exerting downward pressure on restaurant employment there. This mechanism could explain the observed decline in Pennsylvania employment, which drives the positive treatment effect for New Jersey.
If interference is present, it is no longer possible to claim that the increase in the minimum wage led to an increase in employment. Instead, conditional on the parallel trends assumption (Assumption 7) holding, we can only determine the differential effect. That is, we could conclude that the total average effect of the raise in minimum wage on employment in New Jersey was 2.75 FTE employees larger than the average spillover effect in Pennsylvania. However, we cannot determine the effect on employment in any of the states. In fact, it is possible that the rise in minimum wage decreased employment in both states, with a larger decrease in Pennsylvania, resulting in a positive difference but contradicting the original claim in Card and Krueger, (1994).
@@
There are several recent examples published in high-impact journals that do not explicitly invoke SUTVA or the no-interference assumption, but nonetheless discuss potential complications arising from spillover effects. These papers challenge the interpretation of the DiD estimates, characterizing them as ‘general equilibrium’ effects that comprises both indirect and direct effects on the treated (Braghieri et al.,, 2022), in line with our Estimand 1; argue about the direction of the bias coming from the spillovers and how the DiD estimate can be interpreted as a lower bound (Truffa and Wong,, 2025), in line with our results in Proposition 6; or try to disentangle the TATTTATT ($\tau^{1}$) and ASC ($\tau^{0}$) by imputing the missing potential outcome $Y_{i2}(0,\boldsymbol{0},\boldsymbol{0})$ either by finding a control group unaffected by the spillovers (Fenizia and Saggio,, 2024) or by defining an exposure mapping to model interference (Chen et al.,, 2025). By postulating the complete set of potential outcomes and defining causal estimands in the presence of spillovers, this technical note aims to clarify and add structure to such discussions.